two runners start at a certain point simultaneously, going the the same directions, the speed fo one runner is two thirds of the speed of the other runner. if at the end of 3 hours they are 6 miles apart, what is the speed of each runner?
Answer by mananth(16946)

one runner speed = x mph
otehr runner speed = 2x/3
..
difference in speeds = x-2x/3
=x/3

distance = 6 miles
time = 3 hours
6/x= 3
x=2 mph
2x/3 = 4 /3 mph